OpenStudy (anonymous):

can someone help me simplify the integral from x to 1 of (t-1) f(t) dt if x = 0? I'm still stuck on how this =0??

OpenStudy (anonymous):

\[\int\limits_{x}^{1}(t-1) f(t) dt \] if x = 0

OpenStudy (anonymous):

What would this look like when I sub in the 0 for the x? How does it work out?

myininaya (myininaya):

I don't understand the question It would look like this: \[\int\limits_{0}^{1}(t-1)f(t) dt \]

OpenStudy (anonymous):

and then how would you evaluate it??

myininaya (myininaya):

what is f?

OpenStudy (anonymous):

no clue ..is never given....

OpenStudy (anonymous):

this is the right half of the problem y(x) = (x-1)\[\int\limits_{0}^{x} t * f(t) dt + x \int\limits_{x}^{1}(t-1) * f(t) dt\]

myininaya (myininaya):

laurie214 where is this problem located? what book? I don't understand right half of the problem? y(x)=x-1? is y(x)=f(x)?

OpenStudy (anonymous):

and I know how to get the value if x=0 for y(x)....but then

OpenStudy (anonymous):

no book...prof problem...let me type the entire thing again...

myininaya (myininaya):

Are you showing this side equals the other side you mentioned?

OpenStudy (anonymous):

\[I first wrote the y(x) original problem....should have written y('x)=\int\limits_{0}^{x} t * f(t) dt + \int\limits_{x}^{1}\]

myininaya (myininaya):

i'm sleepy...i will look at this later... @foolformath you want to look at this?

OpenStudy (anonymous):

cut me off again....\[y'(x)=\int\limits_{0}^{x}t *f(t) + \int\limits_{x}^{1}(t-1) f(t) dt\]

myininaya (myininaya):

Ok so what are we trying to do? Find y(x)?

OpenStudy (anonymous):

and I need to evaluate if x = 0

myininaya (myininaya):

y'(0) or y(0)?

OpenStudy (anonymous):

I know the first part of the problem would result in a 0 since .....this is what I got when I found y"(x)...now I need to evaluate this if x=0 and x=1

OpenStudy (anonymous):

y'(x) , I mean....so looking for y'(0)

OpenStudy (anonymous):

the first part is \[\int\limits_{0}^{0}t * f(t) d\] and would =0 because of the integral from 0 to 0...

OpenStudy (anonymous):

but the second half has me stuck....the \[+ \int\limits_{0}^{1}(t-1) f(t) dt\]

OpenStudy (anonymous):

would this half also = 0?

OpenStudy (anonymous):

bmp had helped check over my work yesterday when I was finding the first and second derivatives of the original y(x) expression, but he hasn't been available today when I;ve checked in....he told me to send a message, but not sure how to without his name to click on...

myininaya (myininaya):

\[y'(x)=\int\limits_{0}^{x}t f(t) dt+\int\limits_{x}^{1}(t-1) f(t) dt\] So you did get the following: \[y''(x)=xf(x)-(x-1)f(x)=f(x)\]

myininaya (myininaya):

\[y''(x)=f(x)\]

myininaya (myininaya):

\[y'(x)=\int\limits_{}^{}f(x) dx +C\]

OpenStudy (anonymous):

yes :-) but now I have to evaluate y(0), y(1), y'(0) and y'(1)

myininaya (myininaya):

\[\int\limits_{}^{}f(x)dx+C=\int\limits_{0}^{x}t f(t) dt+\int\limits_{x}^{1}(t-1)f(t) dt\]

myininaya (myininaya):

Nothing else is given?

myininaya (myininaya):

No conditions?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

just f is integrable on [0,1}

OpenStudy (anonymous):

how do get the value of an integral that has x=0 yet the terms to be integrated are in terms of t?

myininaya (myininaya):

\[\int\limits_{0}^{x}f(t) dt+C=\int\limits_{0}^{x}tf(t)+\int\limits_{x}^{1}(t-1)f(t) dt \]

myininaya (myininaya):

the limits contain x

myininaya (myininaya):

This is a function of x not t

OpenStudy (anonymous):

ok

myininaya (myininaya):

You want to see an example?

OpenStudy (anonymous):

please! :-)

myininaya (myininaya):

Pretend like \[f(t)=t^2\] \[y'(x)=\int\limits_0^x t \cdot t^2 dt+\int\limits_{x}^{1}(t-1)t^2 dt \] \[y'(x)=\frac{x^4}{4}+\frac{1}{4}-\frac{x^4}{4}-\frac{1}{3}+\frac{x^3}{3}=\frac{x^3}{3}+\frac{-1}{12}=\frac{x^3}{3}+C\] But see for unknown function f(t) I'm trying to figure out that constant :(

myininaya (myininaya):

But see how it is a function of x not t?

myininaya (myininaya):

\[y'(x)=\int\limits_{}^{}f(x) dx+C\]

myininaya (myininaya):

\[y'(x)=\int\limits_{0}^{x}f(t) dt +C\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the value of this would be......?

OpenStudy (anonymous):

if x=0 then this = 0 because the integral is from 0 to 0, or like a to a....

OpenStudy (anonymous):

I get that part, but it's the last part \[\int\limits_{0}^{1}(t-1)f(t) dt\] I'm not sure of

myininaya (myininaya):

\[\int\limits_{0}^{x}f(t) dt+C=\int\limits_{0}^{x}tf(t) dt+\int\limits_{x}^{1}(1-t)f(t) dt\] So if x=0 we have \[\int\limits_{0}^{0}f(t) dt +C=\int\limits_{0}^{0}tf(t) dt+\int\limits_{0}^{1}(1-t)f(t) dt \] \[C=\int\limits_{0}^{1}(1-t)f(t) dt \] But I don't see how this helps :(

myininaya (myininaya):

\[\int\limits_{0}^{x}f(t)dx+\int\limits_{0}^{1}(1-t) f(t) dt=\int\limits_{0}^{x}tf(t)+\int\limits_{x}^{1}(1-t)f(t) dt \]

OpenStudy (anonymous):

I know...I can't get any further...I think he said the value of y' when x =1 was a 0, but that would make the last part a 0 also, and I'm not sure how that happens...

OpenStudy (anonymous):

that's a (t-1) in the parentheses

myininaya (myininaya):

So if x=1 we have \[\int\limits_{0}^{1}f(t) dt+\int\limits_{0}^{1}(1-t) f(t) dt=\int\limits_{0}^{1}tf(t) dt +\int\limits_{1}^{1}(1-t)f(t) dt \] So this gives us \[\int\limits_{0}^{1}f(t) dt +\int\limits_{0}^{1}f(t) dt -\int\limits_{0}^{1}tf(t) dt =\int\limits_{0}^{1}t f(t) dt +0\] this means we have \[2\int\limits_{0}^{1}f(t) dt-\int\limits_{0}^{1}t d(t) dt=\int\limits_{0}^{1}tf(t) dt \] did I do something wrong?

myininaya (myininaya):

Because this means we have \[2\int\limits_{0}^{1} f(t) dt=2 \int\limits_{0}^{1}t f(t) dt \]

myininaya (myininaya):

which means \[\int\limits_{0}^{1}f(t) dt=\int\limits_{0}^{1}t f(t) dt \]

OpenStudy (anonymous):

I have what you have down to the next to last line...

OpenStudy (anonymous):

now how did you get the 2 integral etc.....on the last line ?

myininaya (myininaya):

I wonder if this is true for our little example lets see what happens for \[f(t)=t^2\]

myininaya (myininaya):

i added on both side int(0 to 1 of tf(t) dt)

myininaya (myininaya):

\[\int\limits_{0}^{1}t^2 dt =? \int\limits_{0}^{1}t\cdot t^2 dt \] I'm pretty sure this is not true :(

myininaya (myininaya):

well actually i know it isn't true lol 1/3 does not equal 1/4

myininaya (myininaya):

@bmp I will mention bmp here but you can also do it the message thing laurie like act as if you are going to message me but instead delete my name and put bmp ok?

myininaya (myininaya):

i need sleep laurie I'm sorry It is way passed my bed time

OpenStudy (anonymous):

ok, thanks a bunch! I need sleep, too...tried but couldn't ..this was still on my mind! It's 5:30 am here....awake all night.. :-(

myininaya (myininaya):

omg lol its 4:35 AM here

OpenStudy (anonymous):

we both need sleep! :-) thanks again!

myininaya (myininaya):

i think with a fresher brain we both can work better at this problem well hopefully lol

myininaya (myininaya):

i will see you later today maybe

myininaya (myininaya):

@estudier

OpenStudy (anonymous):

I sent bmp a message...thanks!

OpenStudy (anonymous):

hey can you take a pic of the problem and put it up it will be easier

OpenStudy (anonymous):

This question is not clear at all.

myininaya (myininaya):

lol one more question did you say y(x)=x-1 if so then y'(x)=1 and y''(x)=0 but we had y''(x)=f(x) So f(x)=0 ?

myininaya (myininaya):

ok now i go to sleep

OpenStudy (anonymous):

@myininaya you made a mistake when you chose the limits its true that y''(x)=f(x) but the limits you chose for integrating should be x to 1 and not 0 to x.\[\int\limits_{0}^{x} tf(t)dt+\int\limits_{x}^{1} tf(t)dt+\int\limits_{x}^{1}f(t)dt =\int\limits_{0}^{1}tf(t)dt+\int\limits_{x}^{1}f(t)dt\]

OpenStudy (anonymous):

so \[C=\int\limits_{0}^{1}tf(t)dt. \] which gets cancelled off during the second differentiation.

OpenStudy (anonymous):

so we definately need the function f(t) to go any further we are back at where we started